# TA-KAND: Two-stage Attention Triple Enhancement and U-KAN based Diffusion For Few-shot Knowledge Graph Completion

### v1

@@ 4.2 Result and Comparison With Variants
The report in Table 2 shows that our framework performs better on both datasets across all metrics. Compared to the highest metric results in the aforementioned models, TA-KAND achieves relative improvements of 54% / 53% / 58% / 54% in MRR / Hits@10 / Hits@5 / Hits@1 on NELL, and relative improvements of 34% / 10%/ 19% / 39% on Wiki.
Fig. 2: Visualization of entity-pair embeddings about relation geopolitical-location-residence of person.
To validate our hypothesis that a triple enhancer that differentially extracts features of positive and negative samples can assist the generative model in producing aligned outcomes, we devise distinct enhancers for controlled experimental comparison. Moreover, we pay attention to the performance of U-KAN based diffusion. The results are displayed in Table 3.
(A) Triple Enhancer Variants: In A1, we use the simple attention-based enhancer, where neighbor embedding is coupled without being learned. In A2, the enhancement of entities are computed as the weighted sum of neighbors up to 2-hop, following NP-FKGC [8]. The results demonstrate that enhancer fully exploits neighbor information will improve the accuracy of tail prediction. Moreover, we use t-SNE to visualize the enhanced pairs, in Fig. 2, which partly explains that, if positive and negative samples are distinguished through our enhancement, diffusion process will model the distribution of positive and negative triple rules accurately, thereby better leveraging contrastive signals to select the optimal tail.
(B) Generative Learner Variants: In B1, we replace diffusion learner with neural process as mentioned in NP-FKGC [8] to generate latent semantic representation. In B2, we employ UNet-based diffusion. Experiments show that, given a strong encoder, both neural process and diffusion model can effectively learn latent triple rules. Moreover, inserting KAN layers into a U-Net architecture proves highly compatible and boosts non-linear fitting capacity.

### v2

@@ 4.2 Result and Comparison With Variants
The report in Table 1 shows that our framework performs better on both datasets across all metrics. Compared to the highest metric results in the aforementioned models, TA-KAND achieves significant improvements.
To validate components in our framework. In V1, we use the simple triple enhancer, where neighbor embedding is coupled without being learned. In V2, we replace diffusion learner with neural process as mentioned in NP-FKGC [8] to generate latent semantic representation. In V3, we employ UNet-based diffusion. Experiments show that, given a strong encoder, both neural process and diffusion model can effectively learn latent triple rules. Moreover, inserting KAN layers into a U-Net architecture proves highly compatible and boosts non-linear fitting capacity.